OpenStudy (anonymous):

Activity title: solving problems for uniformly accelerated rectilinear motion learning target: to calculate displacement,velocity,acceleration and time using the four kinematic equations. 1. A french high-velocity train accelerated uniformly from 250km/hr to 300 km/hr as it sped through a distance of 50 km.finds its acceleration.

OpenStudy (anonymous):

i told u about it

OpenStudy (anonymous):

didn't i?

OpenStudy (anonymous):

ishann can i ask u something about his answer

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

v² - u² ------ = a 2 s ( 250 )^2 ( 625 )^2 m^2 (------ ) - (------) ------- ( 3 ) ( 9 ) s^2 ----------------------------------- = a 2 * 50000 m 62500 390625 m^2 ------ - -------- ------ 9 81 s^2 -------------------------- = a 100000 m 562,500 - 390,625 m^2 ------------------ ------ 81 s^2 ------------------------- = a 100,000 m 171,875 1 m^2 ------- * ---------- --------- = a 81 100,000 s^2 * m

OpenStudy (anonymous):

how did he get 250 /3 m/s?

OpenStudy (anonymous):

v² = u² + 2as where v = final velocity u = initial velocity s = distance here v = 300 km/h = 300,000 m/3600 s = 250/3 m/s u = 250 km/h = 250,000 m/3600s = 625/9 m/s s = 50 km = 50000 m he said like this but how did he get 250/3 m/s ?